find the equation of the line through (1, -6) parallel to x+2y=6
Answer by Cromlix(4381)

Change x + 2y = 6 into y = mx + c form
= 2y = -x + 6
Lines that are parallel have the same
gradient/ slope/m value
m1 = m2
So m = - 1
using equation: y -b = m(x - a)
y - (-6) = -1(x - 1)
y + 6 = -x + 1
y = -x + 1 - 6
y = -x - 5.
Hope this helps.
:-)